OpenStudy (anonymous):

solve for x: ln(8-8x)=4

OpenStudy (anonymous):

You have to use euler in both sides in order to take out "ln"

OpenStudy (anonymous):

You can do the rest ; )

OpenStudy (anonymous):

ln(8-8x)=4, 8-8x=e^4, x=(e^4-8)/(-8)=-1/8*e^4+1.